OpenStudy (anonymous):

Solve: 1/9=27^2x-2 dont get it! pls help!...

OpenStudy (dumbcow):

\[\frac{1}{9} = 3^{-2}\] \[27 = 3^{3}\] \[\rightarrow 3^{-2} = 3^{3(2x-2)}\] set exponents equal \[3(2x-2) = -2\] solve for x

OpenStudy (anonymous):

thank you

OpenStudy (dumbcow):

welcome :) btw you should get x=2/3

OpenStudy (anonymous):

yes tats wat i got